OpenStudy (anonymous):

How exactly can I change this to slope intercept method ? y - 5 = 2/3 ( x - 4 ) ?

OpenStudy (anonymous):

y=2/3x+7/3

OpenStudy (anonymous):

first you distribute the 2/3 to the x and -4 then your plus 5 on each side to get the final answer

OpenStudy (anonymous):

I got that but wasn't sure if it was right because i have to graph this, i am not exactly sure how to graph that, so i thought i was wrong.

OpenStudy (anonymous):

Does anybody know how i can graph that ?

OpenStudy (anonymous):

slope intercept form is y = mx + b. y - 5 = 2/3 (x - 4) First thing you have to do is get rid of the fraction. You can do this by multiplying everything by 3. (3) y - 5 = 2/3 (x - 4) 3y - 15 = 2 (x - 4) now multiply through the parenthesis 3y - 15 = 2x - 8 (You want 3y by itself so you add 15 to both sides) 3y - 15 + 15 = 2x - 8 + 15 3y = 2x - 8 + 15 3y = 2x + 7 (however you want y by itself so divide by 3) 3y/3 = 2x/3 + 7/3 y = 2/3x + 7/3

OpenStudy (anonymous):

Thank you Soo much !! Hey, do you know exactly how to graph that though ? I am confused because

OpenStudy (anonymous):

Rise over run.... http://www.algebra.com/algebra/homework/Graphs/Graphs.faq.question.22174.html